(f) y+y=f(t – 37) cos(t), y(0) = 0, y(0) = 1. (g) y + 2y = U(t – 7) +38(t – 37/2) – Ut – 27), y(0) = y(0) = 0.

solve using Laplace transforms
Answer #1

y+2y= ult-T )+38(+-31) - uct-20) L(1)+ QLly)= Llult-TT)+ 3L (814-31) - Lluct-2)) SlLly)-Sy(02-4(0)+244)=ē TS 33118/ 2-21TS